# On absolutely continuous spectrum for one-channel unitary operators

### v1

@@ 2.1. One-dimensional quantum walks
$W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\quad\text{and}\quad Q_{n}=I\quad\text{implying}\quad P_{n}=\mathbf{0}.$
@@
$T_{n}^{\flat}=\varphi_{\flat}(W_{n})=I\,\quad\text{and}\quad\Phi_{(n,+)}=\Psi_{(n+1,-)}=\Psi_{n+1,\uparrow}.$
@@ 2.2. Generalized one-channel quantum walks
Within each ${\mathbb{S}}_{n}$ we assign some ’spin up’ and ’spin down’ orbitals, $(n,\uparrow)$, $(n,\downarrow)$ $\in{\mathbb{S}}_{n}$ which are distinct. We also have a coin operator
@@ 2.3. Stroboscopic unitary dynamics on ${\mathbb{Z}}^{2}$
  • clockwise shift for $n\in{\mathbb{Z}}_{+}$ even where
@@
  • counter-clockwise shift for $n\in{\mathbb{Z}}_{+}$ odd where
@@ 3. Transfer matrix and Resolvent
Comments. From the last formulas above, we may focus on the operator ${\mathcal{U}}^{(u)}$ and include the boundary condition at the root by replacing $T_{z,0}$ with $T_{z,0}\begin{pmatrix}1&0\\ 0&u^{-1}\end{pmatrix}$, while focusing on the operator $\tilde{\mathcal{U}}^{(u)}$ one might include the boundary condition by replacing $T_{z,0}$ with $T_{z,0}\begin{pmatrix}u&0\\ 0&1\end{pmatrix}$.
@@
Using Proposition A.2 this leads to
@@
Then the relations (3.5) hold replacing $R$ with $R^{(u,v)}_{z,[0,0]}$ (o $\tilde{R}^{(u,v)}_{z,[0,0]}$) and $T_{z,0}$ with the corresponding matrix as above, meaning
@@
$\begin{pmatrix}1&0\\ 0&v^{-1}\end{pmatrix}T_{z,0}\begin{pmatrix}u&0\\ 0&1\end{pmatrix}\,=\,\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\varphi_{\sharp}(R^{(u,v)}_{z,[0,0]})\begin{pmatrix}1&0\\ -1&1\end{pmatrix},$
@@
Now let us look at the effects of ‘grouping’ the first $N+1$, shells ${\mathbb{S}}_{0}$, …, ${\mathbb{S}}_{N}$ into a single shell ${\mathbb{G}}_{N}$. Then, using the splitting ${\mathbb{G}}={\mathbb{G}}_{N}\sqcup\bigsqcup_{n=N+1}^{\infty}{\mathbb{S}}_{n}$ we find
@@
$\widehat{U}^{(u)}\,:=\,\widehat{\mathcal{V}}\,\widehat{\mathcal{W}}^{(u)}\quad\text{and}\quad\breve{\mathcal{U}}^{(u)}\,=\,\widehat{\mathcal{W}}^{(u)}\breve{\mathcal{V}}.$
@@
4. Spectral averaging and weak convergence
Let $\mu^{(u,v)}_{N}$ be the spectral measure of ${\mathcal{U}}^{(u,v)}_{N}$ (or $\tilde{\mathcal{U}}^{(u,v)}$) at the vector $e_{(0,-)}$, meaning,
@@ 4. Spectral averaging and weak convergence
for Borel functions $f$ on the unit circle ${\mathbb{U}}(1)$. Note, it is easy to see the second equality using $f(\tilde{\mathcal{U}}^{(u,v)}_{N})=({\mathcal{W}}^{(u)}_{N})^{*}f({\mathcal{U}}^{(u,v)}_{N}){\mathcal{W}}^{(u)}_{N}$ and ${\mathcal{W}}^{(u)}_{N}e_{(0,-)}=ue_{(0,-)}$ with $|u|=1$. Next, we consider an average over the Haar measure for $v\in{\mathbb{U}}(1)$ and define the measure $\mu^{(u)}_{N}$ by
@@
We will denote the value of the Green’s function by $g^{(u,v)}(z)$, meaning
@@
$g^{(u)}_{N}(z):=\int\frac{1}{z^{-1}w-1}\,{\rm d}\mu^{(u)}_{N}(w)\,=\,\frac{1}{2\pi}\int_{0}^{2\pi}g^{(u,e^{i\varphi})}_{N}\,{\rm d}\varphi,\;$
and the Green’s function for the full operator
$g^{(u)}(z):=e_{(0,-)}^{*}\big{(}z^{-1}{\mathcal{U}}^{(u)}-\mathbf{I}\big{)}^{-1}e_{(0,-)}\;=\;\int\frac{1}{z^{-1}w-1}\,{\rm d}\mu^{(u)}(w)\;.$
Now we want to relate $\mu^{(u)}_{N}$ to the transfer matrix. To this end, we will write
$T_{z,[0,N]}\,=\,\begin{pmatrix}A_{z}&B_{z}\\ C_{z}&D_{z}\end{pmatrix}.$ (4.2)
The omit the dependence on $N$ for some of the following computations.
Using Proposition 3.1 we obtain
$g^{(u,v)}_{N}(z)=\frac{-(B_{z}-v^{-1}D_{z})}{(A_{z}-v^{-1}C_{z})u+B_{z}-v^{-1}D_{z}}.$
Now, extending the definitions to $v\in{\mathbb{C}}$, we have
@@
Now for $|u|=1$ and $|z|<1$ fixed, the function is holomorfic in $v$ for $|v|>|z|$, o holomorfic in $v^{-1}$ for $|v^{-1}|<|z|^{-1}$ and $|z|^{-1}>1$. Thus, the average over ${\mathbb{U}}(1)$ simply means to replace $v^{-1}$ by 00. Thus,
$g^{(u)}_{N}(z)\,=\,\frac{-B_{z}}{A_{z}u+B_{z}}\;.$
Now, consider the Poisson transform of the measure $\mu^{(u)}_{N}$, that is,
@@
$=\frac{|A_{z}u+B_{z}|^{2}-B_{z}\overline{(A_{z}u+B_{z}})-\overline{B_{z}}(A_{z}u+B_{z})}{|A_{z}u+B_{z}|^{2}}\,=\,\frac{|A_{z}|^{2}-|B_{z}|^{2}}{|A_{z}u+B_{z}|^{2}}\;.$
All these equations of course work only if the transfer matrices up to $N$ are defined, that is if $z\not\in{\mathcal{A}}_{N}$. Note for $z=e^{i\varphi}\not\in{\mathcal{A}}_{N}$, we have $T_{z,[0,N]}\in{\mathbb{U}}(1,1)$ which implies $T^{*}_{z,[0,N]}\in{\mathbb{U}}(1,1)$ and as such $|A_{z}|^{2}-|B_{z}|^{2}=1$. Moreover,
$\left(T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}\right)^{*}T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}$ $=\begin{pmatrix}u^{*}&1\end{pmatrix}T^{*}_{z,[0,N]}\begin{pmatrix}1\\ &-1\end{pmatrix}T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}$
@@
Thus, for $e^{i\varphi}\not\in{\mathcal{A}}_{N}$ we find
@@
Let us note that ${\rm d}\mu^{(u)}_{N}(e^{i\varphi})$ is the weak limit of $\frac{1}{2\pi}P^{(u)}(re^{i\varphi})\,{\rm d}\varphi$ for $r\nearrow 1$. Hence, the measure $\mu^{(u)}_{N}$ is absolutely continuous in ${\mathbb{U}}(1)\setminus{\mathcal{A}}_{N}$ with respect to the normalized Haar measure on ${\mathbb{U}}(1)$ and the density is given by the right hand side of (4.3) divided by $2\pi$.
As ${\mathcal{A}}_{N}$ is a finite set, the restriction of $\mu^{(u)}_{N}$ to ${\mathcal{A}}_{N}$ is a point measure, let us call this measure $\nu_{N}^{(u)}$. If for $a\in{\mathcal{A}}_{N}$ we have $\nu_{N}^{(u)}(\{a\})=\mu^{(u)}_{N}(\{a\})>0$, then this means that $a$ is an eigenvalue of ${\mathcal{U}}^{(u,v)}_{N}$ for a set of positive measure in $v\in{\mathbb{U}}(1)$. By rank one perturbation arguments, the spectral average over $v$ gives the Haar measure at the vector $e_{(N,+)}$, see for instance [10]. Therefore, if $\nu_{N}^{(u)}(\{a\})>0$ then there is an eigenvector $\Psi\in{\mathbb{C}}^{{\mathbb{G}}_{N}}$ in the orthogonal complement of the cyclic space of $e_{(N,+)}$, in particular $\Psi^{*}e_{(N,+)}=0$. Hence, $\Psi$ is an eigenvector and $a$ the corresponding eigenvalue for all $v\in{\mathbb{U}}(1)$ and furthermore, for $M>N$, $\Psi\oplus\mathbf{0}=\hat{\Psi}$ is an eigenfunction of ${\mathcal{U}}^{(u,v)}_{M}$, implying $\nu_{M}(\{a\})\geq\nu_{N}(\{a\})$. Together with the uniform boundedness of the probability measures $\mu^{(u)}_{N}$ this means that there is a limit point measure
$\nu^{(u)}\,=\,\lim_{N\to\infty}\nu_{N}^{(u)}\;,$
@@
which is supported on the set ${\mathcal{A}}$. Now, further note that in fact $\Psi\oplus\mathbf{0}\in\ell^{2}({\mathbb{G}})=\ell^{2}({\mathbb{G}}_{N})\oplus\ell^{2}({\mathbb{G}}\setminus{\mathbb{G}}_{N})$ is also an eigenfunction of ${\mathcal{U}}^{(u)}$ with eigenvalue $a$ in this case. Any finitely supported eigenfunction of ${\mathcal{U}}^{(u)}$ is of this form and induces an eigenfunction of ${\mathcal{U}}^{(u)}_{N}$ for $N$ large enough. Thus, $\nu^{(u)}$ is precisely the part of the spectral measure $\mu^{(u)}$ coming from the space ${\mathbb{H}}^{(u)}_{c}$ generated by finitely supported eigenfunctions and coincides with the measure defined in (1.17).
Finally, using resolvent convergence we have $g^{(u,v)}_{N}(z)\to g^{(u)}(z)$ for any $|z|<1$ as $N\to\infty$. This implies the convergence of the Poisson transforms and hence $\mu^{(u)}_{N}$ converges weakly to $\mu^{(u)}$. Thus, in sumary we find
${\rm d}\mu^{(u)}(e^{i\varphi})\,=\,{\rm d}\nu^{(u)}(e^{i\varphi})\,+\,\lim_{n\to\infty}\frac{1}{\pi}\,\frac{{\rm d}\varphi}{\left\|T_{e^{i\varphi},[0,N]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\right\|^{2}}\;,$
proving Theorem 1.
The proof of Theorem 2 now works exactly the same as in [39, 47, 48, 49, 51]. As for $\varphi\in{\mathbb{R}}$ we find $T_{e^{i\varphi},[0,n]}\in{\mathbb{U}}(1,1)$ , one has
@@
$1\,\leq\,\left\|\begin{pmatrix}1\\ &-1\end{pmatrix}T_{e^{i\varphi},[0,n]}\begin{pmatrix}0\\ -1\end{pmatrix}\right\|\;\left\|T_{e^{i\varphi},[0,n]}\begin{pmatrix}u\\ 1\end{pmatrix}\right\|\,\leq\,\|T_{e^{i\varphi},[0,n]}\|\,\left\|T_{e^{i\varphi},[0,n]}\begin{pmatrix}u\\ 1\end{pmatrix}\right\|$
@@
means that $\varphi\mapsto\frac{1}{\left\|T_{e^{i\varphi},[0,n]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\right\|^{2}}$ is a bounded sequence in $L^{p}(\varphi_{0},\varphi_{1})$ and has a weakly convergent subsequence with a limit function $\varphi\mapsto g(e^{i\varphi})$. Then, Theorem 1 assures that for $f$ cotinuous on ${\mathbb{U}}(1)$ and compactly supported on $\{e^{i\varphi}\,:\,\varphi\in[a,b]\}$ with $\varphi_{0}<a<b<\varphi_{1}$, $b-a\leq 2\pi$ we find
@@
meaning that $\mu^{(u)}-\nu^{(u)}$ is an absolutely continuous measure w.r.t. ${\rm d}\varphi$ on $e^{i(\varphi_{0},\varphi_{1})}$ and an $L^{p}$ density. This shows the first part of Theorem 2. For the second part, let $[a,b]\subset(\varphi_{0},\varphi_{1})$, $f\geq 0$ and $f(e^{i\varphi})>\varepsilon>0$ for all $\varphi\in[a,b]$. Using Jensen’s inequality for the convex function $F(x)=x^{-1/p}$ and the ‘random variable’ $X(z)=\|T_{z,[0,n]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\|^{2p}$ we find

### v2

@@ 2.1. One-dimensional quantum walks
$W_{n}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\quad\text{and}\quad Q_{n}=\mathbf{I}\quad\text{implying}\quad P_{n}=\mathbf{0}.$
@@
$T_{n}^{\flat}=\varphi_{\flat}(W_{n})=\mathbf{I}\,\quad\text{and}\quad\Phi_{(n,+)}=\Psi_{(n+1,-)}=\Psi_{n+1,\uparrow}.$
@@ 2.2. Generalized one-channel quantum walks
Within each ${\mathbb{S}}_{n}$ we assign some ‘spin up’ and ‘spin down’ orbitals, $(n,\uparrow)$, $(n,\downarrow)$ $\in{\mathbb{S}}_{n}$ which are distinct. Note that, despite using notaitons ‘spin up’ and ‘spin down’, ${\mathbb{S}}_{n}$ may have many more orbitals. In order to define the quantum walk we just pick two for each $n$. As we want to show the analogue structure with the quantum walk and essentially use the same definition for the operators ${\mathcal{S}}$ and ${\mathcal{W}}$, we use the same notation here.
We have a coin operator
@@ 2.3. Stroboscopic unitary dynamics on ${\mathbb{Z}}^{2}$
  • clockwise shift for $n\in 2{\mathbb{Z}}_{+}$ ($n$ even) where
@@
  • counter-clockwise shift for $n\in 2{\mathbb{Z}}_{+}+1$ ($n$ odd) where
@@ 3. Transfer matrix and Resolvent
Remark 3.2.
From the last formulas above, we may focus on the operator ${\mathcal{U}}^{(u)}$ and include the boundary condition at the root by replacing $T_{z,0}$ with $T_{z,0}\begin{pmatrix}1&0\\ 0&u^{-1}\end{pmatrix}$, while focusing on the operator $\tilde{\mathcal{U}}^{(u)}$ one might include the boundary condition by replacing $T_{z,0}$ with $T_{z,0}\begin{pmatrix}u&0\\ 0&1\end{pmatrix}$.
@@
Using Proposition A.2 b) this leads to
@@
Then the relations (3.5) hold replacing $R$ with $R^{(u,v)}_{z,[0,0]}$ (or $\tilde{R}^{(u,v)}_{z,[0,0]}$) and $T_{z,0}$ with the corresponding matrix as above, meaning
@@
$\begin{pmatrix}1&0\\ 0&v^{-1}\end{pmatrix}T_{z,0}\begin{pmatrix}u&0\\ 0&1\end{pmatrix}\,=\,\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\varphi_{\sharp}(\tilde{R}^{(u,v)}_{z,[0,0]})\begin{pmatrix}1&0\\ -1&1\end{pmatrix},$
@@
Now, let us look at the effects of ‘grouping’ the first $N+1$ shells ${\mathbb{S}}_{0}$, …, ${\mathbb{S}}_{N}$ into a single shell ${\mathbb{G}}_{N}$. Then, using the splitting ${\mathbb{G}}={\mathbb{G}}_{N}\sqcup\bigsqcup_{n=N+1}^{\infty}{\mathbb{S}}_{n}$ we find
@@
$\widehat{{\mathcal{U}}}^{(u)}\,:=\,\widehat{\mathcal{V}}\,\widehat{\mathcal{W}}^{(u)}\quad\text{and}\quad\breve{\mathcal{U}}^{(u)}\,=\,\widehat{\mathcal{W}}^{(u)}\breve{\mathcal{V}}.$
@@
4. Proof of Theorem 1 and Theorem 2
In this section we will first prove Theorem 1 using spectral averaging techniques, and then obtain Theorem 2 as a corollary. Let $\mu^{(u,v)}_{N}$ be the spectral measure of ${\mathcal{U}}^{(u,v)}_{N}$ (or $\tilde{\mathcal{U}}^{(u,v)}$) at the vector $e_{(0,-)}$, meaning,
@@ 4. Proof of Theorem 1 and Theorem 2
for Borel functions $f$ on the unit circle ${\mathbb{U}}(1)$. Note, the second equality is easy to see using $f(\tilde{\mathcal{U}}^{(u,v)}_{N})=({\mathcal{W}}^{(u)}_{N})^{*}f({\mathcal{U}}^{(u,v)}_{N}){\mathcal{W}}^{(u)}_{N}$ and ${\mathcal{W}}^{(u)}_{N}e_{(0,-)}=ue_{(0,-)}$ with $|u|=1$. Next, we consider an average over the Haar measure for $v\in{\mathbb{U}}(1)$ and define the measure $\mu^{(u)}_{N}$ by
@@
We will denote the value of the Green’s function by $g_{N}^{(u,v)}(z)$, meaning
@@
$g^{(u)}_{N}(z):=\int\frac{1}{z^{-1}w-1}\,{\rm d}\mu^{(u)}_{N}(w)\,=\,\frac{1}{2\pi}\int_{0}^{2\pi}g^{(u,e^{i\varphi})}_{N}(z)\,{\rm d}\varphi\;.$
Next, we want to relate $\mu^{(u)}_{N}$ to the transfer matrix.
Lemma 4.1.
Denote the entries of the transfer matrix by
$T_{z,[0,N]}\,=\,\begin{pmatrix}A_{z}&B_{z}\\ C_{z}&D_{z}\end{pmatrix}.$
Then, for $|z|<1$, $z\not\in\widehat{\mathcal{A}}_{N}$, the averaged Green’s function is given by
$g^{(u)}_{N}(z)\,=\,\frac{-B_{z}}{A_{z}u+B_{z}}\;.$
Proof.
Using Proposition 3.1 one has
$\varphi_{\sharp}(R_{z,[0,N]}^{(u,v)})$ $=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}\begin{pmatrix}vA_{z}&vB_{z}u^{-1}\\ C_{z}&D_{z}u^{-1}\end{pmatrix}\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$
$=\begin{pmatrix}vA_{z}-C_{z}+vB_{z}u^{-1}-D_{z}u^{-1}&vB_{z}u^{-1}-D_{z}u^{-1}\\ C_{z}+D_{z}u^{-1}&D_{z}u^{-1}\end{pmatrix}.$
Note that if $R_{z,[0,N]}^{(u,v)}=\left(\begin{smallmatrix}\alpha&\beta\\ \gamma&\delta\end{smallmatrix}\right)$, then $g^{(u,v)}_{N}(z)=\alpha$ is the upper right entry and $\varphi_{\sharp}(R_{z,[0,N]}^{(u,v)})\,=\,\left(\begin{smallmatrix}\beta^{-1}&-\beta^{-1}\alpha\\ \delta\beta^{-1}&\gamma-\delta\beta^{-1}\alpha\end{smallmatrix}\right)$. Thus, using $\alpha=\beta^{-1}\alpha/\beta^{-1}$ we obtain
$g^{(u,v)}_{N}(z)=\frac{-(vB_{z}u^{-1}-D_{z}u^{-1})}{vA_{z}-C_{z}+vB_{z}u^{-1}-D_{z}u^{-1}}=\frac{-(B_{z}-v^{-1}D_{z})}{(A_{z}-v^{-1}C_{z})u+B_{z}-v^{-1}D_{z}}.$
Extending the definitions to $v\in{\mathbb{C}}$, we have
@@
Now for $|u|=1$ and $|z|<1$ fixed, the function is holomorphic in $v$ for $|v|>|z|$, or holomorphic in $v^{-1}$ for $|v^{-1}|<|z|^{-1}$ where $|z|^{-1}>1$. Thus, the average over $v\in{\mathbb{U}}(1)$ simply means to replace $v^{-1}$ by 00, giving the desired formula. ∎
Lemma 4.2.
There is a point measure $\nu_{N}^{(u)}$ supported on the finite set ${\mathcal{A}}_{N}$, such that for $f\in C({\mathbb{U}}(1))$
$\mu^{(u)}_{N}(f)\,=\,\nu^{(u)}_{N}(f)\,+\,\int_{0}^{2\pi}\frac{f(e^{i\varphi})}{\pi\|T_{e^{i\varphi},[0,N]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\|^{2}}\,{\rm d}\varphi\,.$
Proof.
Consider the Poisson transform of the measure $\mu^{(u)}_{N}$, that is, using Lemma 4.1
@@
$=\frac{|A_{z}u+B_{z}|^{2}-B_{z}\overline{(A_{z}u+B_{z}})-\overline{B_{z}}(A_{z}u+B_{z})}{|A_{z}u+B_{z}|^{2}}\,=\,\frac{|A_{z}|^{2}-|B_{z}|^{2}}{|A_{z}u+B_{z}|^{2}}\;$
for $|z|<1$, $z\not\in\widehat{\mathcal{A}}_{N}$. Note, for $z=e^{i\varphi}\not\in{\mathcal{A}}_{N}=\widehat{\mathcal{A}}_{N}\cap{\mathbb{U}}(1)$ we have $T_{z,[0,N]}\in{\mathbb{U}}(1,1)$ which implies $T^{*}_{z,[0,N]}\in{\mathbb{U}}(1,1)$ and as such $|A_{z}|^{2}-|B_{z}|^{2}=1$. Moreover,
$\left(T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}\right)^{*}\begin{pmatrix}1\\ &-1\end{pmatrix}T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}$ $=\begin{pmatrix}u^{*}&1\end{pmatrix}T^{*}_{z,[0,N]}\begin{pmatrix}1\\ &-1\end{pmatrix}T_{z,[0,N]}\begin{pmatrix}u\\ 1\end{pmatrix}$
@@
Thus, for $e^{i\varphi}\not\in{\mathcal{A}}_{N}$
@@
Let us note that ${\rm d}\mu^{(u)}_{N}(e^{i\varphi})$ is the weak limit of $\frac{1}{2\pi}P^{(u)}(re^{i\varphi})\,{\rm d}\varphi$ for $r\nearrow 1$. Hence, the measure $\mu^{(u)}_{N}$ is absolutely continuous in ${\mathbb{U}}(1)\setminus{\mathcal{A}}_{N}$ with respect to the normalized Haar measure on ${\mathbb{U}}(1)$ and the density is given by the right hand side of (4.2) divided by $2\pi$. Moreover, as ${\mathcal{A}}_{N}$ is a finite set, the restriction of $\mu^{(u)}_{N}$ to ${\mathcal{A}}_{N}$ is a point measure $\nu_{N}^{(u)}$. This finishes the proof. ∎
Lemma 4.3.
We have $\nu^{(u)}_{N+1}\geq\nu^{(u)}_{N}$ and $\lim_{N\to\infty}\nu_{N}^{(u)}=\nu^{(u)}$ with $\nu^{(u)}$ as defined in (1.17).
Proof.
If for $a\in{\mathcal{A}}_{N}$ we have $\nu_{N}^{(u)}(\{a\})=\mu^{(u)}_{N}(\{a\})>0$, then this means that $a$ is an eigenvalue of ${\mathcal{U}}^{(u,v)}_{N}$ for a set of positive measure in $v\in{\mathbb{U}}(1)$. By rank one perturbation arguments, at the vector $e_{(N,+)}$, the spectral average over $v=e^{i\varphi}\in{\mathbb{U}}(1)$ with respect to the Haar measure gives the Haar measure on ${\mathbb{U}}(1)$, see for instance [10, Proposition 8.1]333where in the notations of [10] we use Proposition 8.1 with $T=1$, $e^{-iH_{0}}={\mathcal{U}}_{N}^{(u,1)}$, $\phi=e_{(N,+)}$ and the average over $\kappa\in{\mathbb{T}}={\mathbb{R}}/2\pi{\mathbb{Z}}$ corresponds to the average over $\varphi$. This means, for $f\in C({\mathbb{U}}(1))$
$\int_{0}^{2\pi}\,\left\langle e_{(N,+)},f\left({\mathcal{U}}^{(u,e^{i\varphi})}_{N}\right)e_{(N,+)}\,\right\rangle\frac{\rm d\varphi}{2\pi}$
$\quad=\int_{0}^{2\pi}\,\left\langle e_{(N,+)},f\left(e^{i\varphi|e_{(N,+)}\rangle\langle e_{(N,+)}|}\,{\mathcal{U}}^{(u,1)}_{N}\right)e_{(N,+)}\,\right\rangle\frac{\rm d\varphi}{2\pi}=\int_{0}^{2\pi}f(e^{i\theta})\frac{\rm d\theta}{2\pi}\,.$ (4.3)
Now fix $a\in{\mathbb{U}}(1)$ and consider the restriction of ${\mathcal{U}}^{(u,v)}_{N}$ to the cyclic space ${\mathcal{Z}}_{N}$ generated by $e_{(N,+)}$. Note that ${\mathcal{Z}}_{N}$ does not depend on $v$. By (4.3), the set
$\{v\in{\mathbb{U}}(1)\,:\,a\,\text{is eigenvalue of the restriction }{\mathcal{U}}^{(u,v)}_{N}|_{{\mathcal{Z}}_{N}}\,\}$
has Haar measure zero. Therefore, if $\nu_{N}^{(u)}(\{a\})>0$, then, for some $v$ there is an eigenvector $\Psi\in{\mathcal{Z}}_{N}^{\perp}$, meaning $\Psi^{*}e_{(N,+)}=0$. Such an eigenvector $\Psi$ is in fact an eigenvector of ${\mathcal{U}}^{(u,v)}_{N}$ for all $v\in{\mathbb{U}}(1)$, with the same eigenvalue $a$. Thus, we find that $\nu^{(u)}_{N}(\{a\})$ is precisely given by the norm squared of $e_{(0,-)}$ projected to $\ker({\mathcal{U}}^{(u,v)}_{N}-a\mathbf{I})\cap{\mathcal{Z}}_{N}^{\perp}$.
Furthermore, if $\Psi\in\ker({\mathcal{U}}^{(u,v)}_{N}-a\mathbf{I})\cap{\mathcal{Z}}_{N}^{\perp}$, then, for $M>N$ we find that $\Psi\oplus\mathbf{0}=\hat{\Psi}\in\ker({\mathcal{U}}^{(u,v)}_{M}-a\mathbf{I})\cap{\mathcal{Z}}_{M}^{\perp}$, implying, $\nu_{M}(\{a\})\geq\nu_{N}(\{a\})$. Together with the uniform boundedness of the probability measures $\mu^{(u)}_{N}$ this means that there is a limit point measure
$\tilde{\nu}^{(u)}\,=\,\lim_{N\to\infty}\nu_{N}^{(u)}\;,$
@@
which is supported on the set ${\mathcal{A}}$. Further note that, in fact, $\Psi\oplus\mathbf{0}\in\ell^{2}({\mathbb{G}})=\ell^{2}({\mathbb{G}}_{N})\oplus\ell^{2}({\mathbb{G}}\setminus{\mathbb{G}}_{N})$ is also an eigenfunction of ${\mathcal{U}}^{(u)}$ with eigenvalue $a$ in this case. Any finitely supported eigenfunction of ${\mathcal{U}}^{(u)}$ is of this form and induces an eigenfunction of ${\mathcal{U}}^{(u,v)}_{N}$ in ${\mathcal{Z}}_{N}^{\perp}$ for $N$ large enough. Thus, $\tilde{\nu}^{(u)}$ is precisely the part of the spectral measure $\mu^{(u)}$ coming from the space ${\mathbb{H}}^{(u)}_{c}$ generated by finitely supported eigenfunctions and $\tilde{\nu}^{(u)}=\nu^{(u)}$ as defined in (1.17). ∎
Now we are ready to prove Theorem 1.
Proof of Theorem 1.
Using resolvent convergence we have
$g^{(u,v_{N})}_{N}(z)\to g^{(u)}(z):=e_{(0,-)}^{*}\big{(}z^{-1}{\mathcal{U}}^{(u)}-\mathbf{I}\big{)}^{-1}e_{(0,-)}\;=\;\int\frac{1}{z^{-1}w-1}\,{\rm d}\mu^{(u)}(w)\;.$
for any $|z|<1$ and any sequence $v_{N}\in{\mathbb{U}}(1)$ as $N\to\infty$. In particular, this means that the compact sets $\{g^{(u,v)}_{N}(z)\,:\,v\in{\mathbb{U}}(1)\}$ shrink to a point and we also find convergence for the averages, $g^{(u)}_{N}(z)\to g^{(u)}(z)$. This implies the convergence of the Poisson transforms and hence $\mu^{(u)}_{N}$ converges weakly to $\mu^{(u)}$. Thus, with Lemma 4.2 and Lemma 4.3 we find
${\rm d}\mu^{(u)}(e^{i\varphi})\,=\,{\rm d}\nu^{(u)}(e^{i\varphi})\,+\,\lim_{n\to\infty}\frac{1}{\pi}\,\frac{{\rm d}\varphi}{\left\|T_{e^{i\varphi},[0,N]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\right\|^{2}}\;$
weakly, proving Theorem 1. ∎
Proof of Theorem 2.
The proof works exactly the same as in [39, 47, 48, 49, 51]. As for $\varphi\in{\mathbb{R}}$ we find $T_{e^{i\varphi},[0,n]}\in{\mathbb{U}}(1,1)$ , one has
@@
$1\,\leq\,\left\|\begin{pmatrix}1\\ &-1\end{pmatrix}T_{e^{i\varphi},[0,n]}\begin{pmatrix}0\\ -1\end{pmatrix}\right\|\;\left\|T_{e^{i\varphi},[0,n]}\begin{pmatrix}u\\ 1\end{pmatrix}\right\|\,\leq\,\|T_{e^{i\varphi},[0,n]}\|\,\left\|T_{e^{i\varphi},[0,n]}\begin{pmatrix}u\\ 1\end{pmatrix}\right\|,$
@@
means that $\varphi\mapsto\frac{1}{\left\|T_{e^{i\varphi},[0,n]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\right\|^{2}}$ is a bounded sequence in $L^{p}(\varphi_{0},\varphi_{1})$ and has a weakly convergent subsequence with a limit function $\varphi\mapsto g(e^{i\varphi})$ in $L^{p}$ (note $p>1$). Then, Theorem 1 assures that for $f$ continuous on ${\mathbb{U}}(1)$ and compactly supported on $\{e^{i\varphi}\,:\,\varphi\in[a,b]\}$ with $\varphi_{0}<a<b<\varphi_{1}$, $b-a\leq 2\pi$ we find
@@
meaning that $\mu^{(u)}-\nu^{(u)}$ is an absolutely continuous measure w.r.t. ${\rm d}\varphi$ on $e^{i(\varphi_{0},\varphi_{1})}$ and has an $L^{p}$ density. This shows the first part of Theorem 2. For the second part, let $[a,b]\subset(\varphi_{0},\varphi_{1})$, $f\geq 0$ and $f(e^{i\varphi})>\varepsilon>0$ for all $\varphi\in[a,b]$. Using Jensen’s inequality for the convex function $F(x)=x^{-1/p}$ and the ‘random variable’ $X(z)=\|T_{z,[0,n]}\left(\begin{smallmatrix}u\\ 1\end{smallmatrix}\right)\|^{2p}$ we find